Calculate the expected return on stock of Gamma Inc.:

State of the economy Probability of the states Percentage returns
Economic recession 27% -7.3%
Steady economic growth      40% 3.9%
Boom Please calculate it 13.7%
Answer #1

Calculation of expected return:

Probability of boom= 1-0.27-0.40= 0.33

Expected return= (weight of economic*return of economic)+(weight of steady economic*return)+(weight of boom*return)

Expected return= (0.27*-7.3)+(0.40*3.9)+(0.33*13.7)= 4.11

Expected return= 7.11%